OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y greater then or equal to 4x - 2? (0, -4) (0, 0) (0, -3) (0, -5)

OpenStudy (anonymous):

y>=4x-2. Pick any of the point that are offered and plug them in. For example using the first one -4>=4(0)-2........-4>=-2....False Try the next one y>=4x-2.... (0,0) 0>=4(0)-2......0>=-2

OpenStudy (anonymous):

Seems like the second one is True